### v2

# FreeBlend: Advancing Concept Blending with Staged Feedback-Driven Interpolation Diffusion

Abstract
Concept blending is a promising yet underexplored area in generative models. While recent approaches, such as embedding mixing and latent modification based on structural sketches, have been proposed, they often suffer from incompatible semantic information and discrepancies in shape and appearance. In this work, we introduce FreeBlend, an effective, training-free framework designed to address these challenges. To mitigate cross-modal loss and enhance feature detail, we leverage transferred image embeddings as conditional inputs. The framework employs a stepwise increasing interpolation strategy between latents, progressively adjusting the blending ratio to seamlessly integrate auxiliary features. Additionally, we introduce a feedback-driven mechanism that updates the auxiliary latents in reverse order, facilitating global blending and preventing rigid or unnatural outputs. Extensive experiments demonstrate that our method significantly improves both the semantic coherence and visual quality of blended images, yielding compelling and coherent results.
Machine Learning, ICML
Figure 1: We introduce FreeBlend, a novel, training-free approach that effectively blends concepts to generate new objects through feedback interpolation and auxiliary inference. Unlike previous methods such as TEXTUAL (Melzi et al., 2023), which often suffer from category-specific bias, or Composable Diffusion (Liu et al., 2022), which tends to generate outputs with co-occurring subjects, FreeBlend provides a more robust solution. Additionally, methods like MagicMix (Liew et al., 2022) rely on the structural similarity between the reference and blending concepts, often resulting in unstable blending outcomes. In contrast, FreeBlend consistently produces visually coherent and harmonious blends, setting a new benchmark for state-of-the-art blending techniques.
1 Introduction
The development of diffusion models (Ho et al., 2020, 2022) has significantly advanced the field of image generation (Rombach et al., 2022; Esser et al., 2023; Brooks et al., 2023; Guo et al., 2023). These models have exhibited impressive abilities of producing novel and appealing images that align with specific instructions, showcasing their potential and flexibility for controlled image synthesis.
Concept blending, as a novel task in the application of diffusion models, involves blending distinct concepts to create a new one that retains the defining characteristics of its components (Fauconnier & Turner, 1998; Ritchie, 2004). It can be viewed as a way of integrating elements from different domains (e.g., cats and dogs) into a novel and meaningful output and creating new objects, scenes, or alterations that are coherent and creatively synthesized. Although current models can generate realistic combinations of individual concepts (Liu et al., 2022; Chefer et al., 2023), more sophisticated concept blending techniques are needed to ensure that the generated images not only blend features but also maintain their logical consistency and visual appeal.
Early methods of concept blending rely on simple vector weighting (Melzi et al., 2023) or hand-crafted feature combinations (Olearo et al., 2024), which involve adding or reassembling text embeddings. These approaches often lead to inaccurate representations and lack correspondence between visual and semantic features due to cross-modal discrepancies. MagicMix (Liew et al., 2022) interpolates the original class latent into another latent space corresponding to a text prompt. While it creatively introduces latent interpolation for mixing, it struggles with shape mismatch issues and lacks flexibility in visual transformations. ConceptLab (Richardson et al., 2024) utilizes VLMs and latent space manipulation, but the constraints imposed during training the additional module limit the flexibility of its application. ATIH (Xiong et al., 2024), based on MagicMix, injects trainable parameters into the denoising process and enforces similarity constraints to harmonize the fusing of texts and images. However, the limitations of its model structure, similar to those of MagicMix, combined with its inability to address mismatched shapes or semantically irrelevant features, hinder its overall blending performance.
In this paper, we introduce FreeBlend, a novel training-free approach to concept blending. At its core, FreeBlend employs a feedback-driven mechanism with latent interpolation to steer the diffusion denoising process. By interpolating auxiliary latents into the initial blending latent and iterating through feedback-driven interactions, the influence of the auxiliary latents gradually diminishes, guiding the blending latent towards the generation of the final image. This integration of blending and denoising processes leverages the strengths of multiple diffusion models to enhance concept blending performance.
Specifically, FreeBlend consists of three core components: transferred unCLIP (Ramesh et al., 2022) image conditions for Stable Diffusion (Rombach et al., 2022), a stepwise increasing interpolation strategy, and a feedback-driven mechanism of denoising process. Instead of using traditional text-based conditions, we employ images generated from text as conditions to guide the generation process via the unCLIP model. This approach provides more precise visual details compared to the vagueness of text-based conditions, reducing the uncertainty caused by cross-modal differences. In addition to this, we divide the denoising process into three stages: the initialization stage, the blending stage, and the refinement stage. At the initialization stage, the pretrained Stable Diffusion model starts with random noise sampled from a unit Normal distribution, which is then denoised under the guidance of unCLIP image condition. At the blending stage, we add noise to the auxiliary latents derived from the condition images to ensure to be in the same period as the blending latent. The blending and auxiliary processes are all denoised simultaneously. In the final refinement stage, only the unCLIP image condition is used to provide additional information, enabling the model to generate images with greater clarity and finer details.
During the blending stage, we set the proportion of the current timestep relative to the total denoising time and use it to adjust the contribution of the auxiliary latents to a blending latent. As denoising progresses, this coefficient decreases, gradually reducing the influence of the auxiliary latents while focusing more on the blending latents, allowing more intricate details to emerge. The feedback-driven mechanism applies this coefficient uniformly, regulating the influence of the auxiliary latents on the blending latents. During the early blending stage, the auxiliary latents enrich the blending latents with information from the original class, and later, they add subtle details without disrupting the overall structure of the image. In addition to qualitative evaluation, we further incorporate quantitative metrics to rigorously assess the generation process, providing a more comprehensive analysis of the model’s performance on the datasets we have generated.
The main contributions of this paper are three-fold:
  • We introduce a novel feedback-driven latent interpolation diffusion model to tackle the concept blending problem. To the best of our knowledge, this is the first training-free method specifically designed for exhaustive concept blending.
  • Our approach incorporates unCLIP to use images as conditions, along with a stepwise increasing interpolation strategy and a feedback-driven denoising process to effectively blend different concepts.
  • Extensive qualitative and quantitative experiments are conducted on several benchmarks, including our proposed blending effect evaluation metric, CLIP-BS, demonstrating that our method achieves state-of-the-art performance in generating blended concepts.
2 Related Work
Image Editing and Training-free Guidance. Generative models like Stable Diffusion (Rombach et al., 2022) have advanced text-to-image synthesis (Podell et al., 2024; Zhang et al., 2023; Gu et al., 2022) and diffusion models have significantly advanced image editing (Balaji et al., 2022; Couairon et al., 2022; Brooks et al., 2023). DiffEdit (Couairon et al., 2022) and InstructPix2Pix (Brooks et al., 2023) primarily focus on semantic image editing. DiffEdit leverages mask guidance to facilitate intuitive modifications, while InstructPix2Pix enables efficient editing through natural language instructions. Shi et al. (2024) propose DragDiffusion, which improves interactive editing by allowing precise adjustments with handle points and editable regions. In training-free guidance, Feng et al. (2022) introduce structured diffusion guidance to tackle compositional challenges, enhancing the handling of multiple objects. Yu et al. (2023) develop FreeDoM, a training-free method using energy-guided conditional diffusion for dynamic editing with conditions like segmentation maps. Mo et al. (2024) propose FreeControl, offering zero-shot spatial control over pretrained models. These methods enhance the flexibility and precision of diffusion models for effective, retraining-free image editing. Our work builds on this intuition and introduces a novel method for achieving a unique blending style of image editing through training-free guidance.
Concept Composition. The field of diffusion models has seen significant progress in concept composition, where multiple concepts are combined to create complex, multi-faceted images. Liu et al. (2022) introduce a compositional approach that uses multiple models to generate distinct components of an image, effectively addressing challenges related to complex object compositions and their interrelationships. Building on this, Kumari et al. (2023) explore multi-concept customization, enabling the generation of unified, complex images that seamlessly blend multiple concepts. Furthermore, Chefer et al. (2023) propose the Attend-and-Excite method, which improves image accuracy by fine-tuning attention mechanisms, allowing for better semantic alignment between objects in the generated image.
Concept Blending. In contrast, while concept blending holds promise, its practical applications remain more limited and less explored compared to compositional approaches. MagicMix (Liew et al., 2022) addresses semantic mixing while preserving spatial layout, but its use is more restricted due to its shape limitations. Melzi et al. (2023) and Olearo et al. (2024) explore concept blending by manipulating the relationship between different prompts, with Melzi et al. (2023) studying prompt ratios and Olearo et al. (2024) testing various mechanisms for blending text embeddings. Richardson et al. (2024) introduce ConceptLab, using Diffusion Prior models to generate novel concepts within a category through CLIP-based constraints, allowing for the creation of unique hybrid concepts. Xiong et al. (2024) further advance the field with adaptive text-image harmony, merging text and image inputs to generate novel objects while preserving their original layout. While these studies contribute to the domain of concept blending, their overall impact has been more limited compared to concept composition, which has motivated the completion of our work.
3 Methodology
3.1 Preliminaries
The Stable Diffusion (Rombach et al., 2022) model efficiently operates in the latent space, distinguishing it from previous approaches (Ho et al., 2020; Dhariwal & Nichol, 2021). Specifically, an autoencoder consisting of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$ is trained with a reconstruction objective. Given an image $\mathbf{x}$, the encoder $\mathcal{E}$ maps it to a latent $\mathbf{z}$, and the decoder $\mathcal{D}$ reconstructs the image from the latent, i.e., $\tilde{\mathbf{x}}=\mathcal{D}(\mathbf{z})=\mathcal{D}(\mathcal{E}(\mathbf{x}))$. This way, at each timestep $t$, a noisy latent $\mathbf{z}_{t}$ can be obtained. Beyond the routine training scheme, Stable Diffusion devises conditioning mechanisms to control the synthesized image content with $\mathbf{c}$ representing given additional conditions like text prompts, images, or other measurable properties. The DDPM (Ho et al., 2020) model $\epsilon_{\theta}$ can then be trained via:
$\mathcal{L}=\mathbb{E}_{\mathbf{z}\sim\mathcal{E}(\mathbf{x}),\mathbf{c},\epsilon\sim\mathcal{N}(0,I),t}\left[\left\|\epsilon-\hat{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{c})\right\|_{2}^{2}\right],$ (1)
where U-Net (Ronneberger et al., 2015) enhanced with self-attention and cross-attention layers is adopted as the denoiser. During training, given a noised latent $\mathbf{z}_{t}$ at timestep $t$ and condition $\mathbf{c}$, denoiser is tasked with predicting the noise $\epsilon_{t}$ added to the current latent.
During inference, a latent $\mathbf{z}_{T}$ is sampled from the standard normal distribution $\mathcal{N}(0,I)$ and the DDPM is used to iteratively remove the noise $\epsilon_{t}=\hat{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{c})$ in $\mathbf{z}_{T}$ to produce $\mathbf{z}_{0}$. This process is expressed as:
$\hat{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{c})\approx-\sigma_{t}\nabla_{\mathbf{z}_{t}}\log p_{t}(\mathbf{z}_{t}|\mathbf{c}),$ (2)
where $\hat{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{c})$ represents the noise removal at timestep $t$, and the gradient of the log marginal distribution provides the direction for noise reduction. In the end, the latent $\mathbf{z}_{0}$ is passed to the decoder $\mathcal{D}$ to generate the output image $\tilde{\mathbf{x}}$.
Figure 2: Overview of our method at the blending stage. Two input images or concepts are encoded into embeddings, mapped to a shared text space via the Linear Prior Converter from unCLIP (Ramesh et al., 2022). These embeddings condition the U-Net: one for downsampling, the other for upsampling. During the blending stage, a blending latent $L_{b}$ initialized with Gaussian noise is processed in the Feedback Interpolation Module, conditioned on image embeddings. Noise $\epsilon$ is added to the embeddings to generate initial auxiliary latents, which are interpolated into $L^{(t)}_{b}$ with an increasing weight $p$. The $L^{(t)}_{a}$ is combined with interpolated latent $L^{\prime(t)}_{b}$ by proportion $p$. All updated $L^{\prime(t)}_{a}$ are refined in the auxiliary inference to retain original features using the text prompt for corresponding categories, and $L^{\prime(t)}_{b}$ is denoised via the blending inference. Finally, the refined $L_{b}$ is passed into the VAE decoder to generate the final blending image.
3.2 unCLIP Image Embeddings for Condition
Previous diffusion models rely on a variety of conditions (Zhan et al., 2024), and traditional concept blending approaches typically use transformed text embeddings (Melzi et al., 2023; Olearo et al., 2024) as inputs to these models. In contrast, our approach utilizes the image embeddings as the sole input, leading to significant improvements in both quality and interpretability. Given that the unCLIP (Ramesh et al., 2022) model, which shares similarities with the DALL-E (Ramesh et al., 2021) mechanism, allows images to serve as conditional inputs, we adopt it as the backbone model for our method.
In order to preprocess $\mathbf{c}$ from the given image modality, we introduce a Linear Prior Converter $\tau_{\theta^{\prime}}$ from unCLIP that maps images from image embeddings into the text embedding space for sampling condition. These embeddings are subsequently used as inputs to the intermediate layers of the U-Net through a cross-attention mechanism, which is defined as
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)\cdot V,$ (3)
where
$\left\{\begin{array}[]{l}Q=W_{Q}\cdot\varphi(L^{(t)}_{b})\\ K=W_{K}\cdot\tau_{\theta^{\prime}}(\mathbf{c}_{\text{stage}})\\ V=W_{V}\cdot\tau_{\theta^{\prime}}(\mathbf{c}_{\text{stage}}),\\ \end{array}\right.$ (4)
and $\mathbf{c}_{\text{stage}}$ is defined as:
$\mathbf{c}_{\text{stage}}=\begin{cases}\mathbf{c}_{\text{up}}&\text{(for downsampling)}\\ \mathbf{c}_{\text{down}}&\text{(for upsampling)}\end{cases}.$ (5)
Here, $\varphi(L^{(t)}_{b})\in\mathbb{R}^{N\times d_{\epsilon}}$ represents the intermediate (flattened) feature map produced by the U-Net that implements $\epsilon_{\theta}$, and $W_{V}\in\mathbb{R}^{d\times d_{\epsilon}}$, $W_{Q}\in\mathbb{R}^{d\times d_{\tau}}$, and $W_{K}\in\mathbb{R}^{d\times d_{\tau}}$ are learnable projection matrices (Vaswani, 2017; Jaegle et al., 2021). $\mathbf{c}_{\text{up}}$ and $\mathbf{c}_{\text{down}}$ are concept conditions of $\mathbf{c}_{\text{stage}}$, and they are all encoded from original images.
By utilizing the first embedding during the downsampling phase and the second embedding during upsampling, the model is able to seamlessly integrate hybrid features at each training epoch. This methodology mitigates the risk of sequentially introducing two distinct feature sets across epochs, a process that could lead to feature conflicts due to the inherent differences in their nature. Furthermore, a simple averaging of embeddings may result in the attenuation of critical information, such as fine details or spatial positioning. Leveraging the ability to interface with the text embedding space via unCLIP, the model not only captures the semantic essence of the concept but also affords precise control over visual attributes, including shape and color.
3.3 Staged Denoising Process
To achieve more flexible and precise control over image generation, motivated by Liew et al. (2022); Lin (2024); Ackermann & Li (2022), we divide the generation phase into three stages: the initialization stage, the blending stage, and the refinement stage. And we use two coefficients, $t_{s}$ and $t_{e}$, to denote the start and end timesteps of the blending stage, respectively.
At the initialization stage, the denoising process using Stable Diffusion begins with random noise $\epsilon\sim\mathcal{N}(0,I)$. In this stage, the blending latent is denoised based on the input unCLIP image embeddings, which define the basic layout of the subject and background in the image. Starting with $L^{(T)}_{b}=\epsilon$, the blending latent is iteratively updated according to the following:
$L^{(t-1)}_{b}=$ $\epsilon_{\theta}(L^{(t)}_{b},t,\varnothing)$ (6)
$+w\cdot\left[\epsilon_{\theta}(L^{(t)}_{b},t,\mathbf{c}_{\text{stage}})-\epsilon_{\theta}(L_{b}^{(t)},t,\varnothing)\right],$
where $w$ denotes the classifier-free guidance (Ho & Salimans, 2022) scale, and $\epsilon_{\theta}(L^{(t)}_{b},t,\mathbf{c})$ is the pre-trained U-Net model with conditional input $\mathbf{c}$ and timestep $t$.
At the blending stage, the initially formed latent is interpolated with auxiliary latents to incorporate original features. In this process, we employ a stepwise increasing interpolation strategy along with a feedback-driven mechanism to blend and filter out irrelevant features from the concepts.
At the refinement stage, the general appearance of $L_{b}$ has been completed. The focus now shifts to enhancing the finer details of the latent, making the result more natural and harmonious while eliminating any sense of disjointedness or artificial stitching. Using the same conditions as in the initialization stage, this phase continues to build upon the previous layout, further refining and improving it.
3.4 Stepwise Increasing Interpolation Strategy
In this section, we elaborate on how our method ensures semantic coherence. We interpolate two image latents with the current denoising latents to encode semantic information from both sources together, which shares similar intuition with MagicMix (Liew et al., 2022). The key challenge is to select appropriate blending ratio at each timestep. One issue of using a constant blending ratio is that we may result in blurring and unclear images, dramatically decreasing the generation quality. To address this, for a denoising process with $T$ timesteps, at timestep $t$, we apply a stepwise declining blending ratio $p$, which could be expressed as
$p=1-\frac{t}{T},$ (7)
where $N$ is the number of concepts needed to blend, and the proportion of the $k$-th auxiliary latent $L^{(k)}_{a}$ is
$\lambda=\frac{1-p}{N}.$ (8)
Another challenge is the potential for biased or catastrophic forgetting of content. Specifically, the model should account for the varying influence of different concepts. For instance, when subject concepts such as “animals” are blended with background concepts like “plants”, the model might give more weight to the subject concepts, resulting in generated images with a higher representation of animals.
To address this, we introduce an additional hyper-parameter, denoted as $\gamma$, which regulates the contribution of each image embedding in the latent space. In the case of a blending task involving $N$ images, the interpolation process can be formulated as follows:
$L^{\prime(t)}_{b}=p\cdot L^{(t)}_{b}+\lambda\cdot\sum_{n=1}^{N}\gamma_{n}\cdot L_{a}^{(t,n)},$ (9)
where $\gamma_{n}$ denotes the interpolation strength of $n$-th image, $L^{(t)}_{b}$ refers to the blending latent, and $L_{a}^{(t,n)}$ specifies the $n$-th auxiliary latent at the $t$-th timestep.
Figure 3: The visualized results highlight the superiority of our approach over other methods. We provide a qualitative comparison of our method with three typical concept blending techniques: TEXTUAL (Melzi et al., 2023), Composable Diffusion (Liu et al., 2022) and MagicMix (Liew et al., 2022). For each method and pair, three images are presented. Compared to other methods, which suffer from rigid splicing, discordant compositions, or bias towards one concept, our method excels at seamlessly integrating the features of different concepts into a cohesive new object.
3.5 Feedback-Driven Mechanism
In terms of synthesis quality, setting the time proportion of the blending stage too high can lead to overlapping issues, particularly when the blending stage ends at a high value. This indicates that significant changes should not be introduced during the later stages of denoising, as this could cause unwanted fluctuations and overlapping effects. In previous methods (Liew et al., 2022), the blending stage ends earlier, which limits the ability for further blending and adjustment. To address this issue, we propose a feedback-driven mechanism.
Specifically, $L^{(t_{s},k)}_{a}$ is initialized using the initial auxiliary latent $L^{(0,k)}_{a}$ as follows:
$L^{(t_{s},k)}_{a}=\sqrt{\bar{\alpha}_{t}}L^{(0,k)}_{a}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$ (10)
where $L^{(0,k)}_{a}$ is the embedding encoded from image.
Once $L^{(t)}_{b}$ is interpolated with $L^{(t,k)}_{a}$, the latter should be simultaneously updated using $L^{\prime(t)}_{b}$ as follows:
$L^{\prime(t,k)}_{a}$ $=p\cdot L^{\prime(t)}_{b}+(1-p)\cdot L_{a}^{(t,k)}$ (11)
$=\underbrace{(1-p)\cdot L_{a}^{(t,k)}}_{\text{inherits the original appearance}}$
$+\underbrace{p^{2}\cdot L^{(t)}_{b}}_{\text{integrates into the subject}}$
$+\underbrace{p\cdot(1-p)\cdot\left[\frac{1}{N}\cdot\sum_{n=1}^{N}\gamma_{n}\cdot L_{a}^{(t,n)}\right]}_{\text{maintains balance}}.$
After interpolation, $L_{a}$ should also be denoised as described in Equation 6, with $\mathbf{c}_{\text{stage}}$ being replaced by $\mathbf{c}_{\text{up}}$ or $\mathbf{c}_{\text{down}}$ according to the category of it.
As shown in Equation 11, as the timesteps decrease, $p$ increases, which causes $L_{b}$ to take up a larger proportion of $L_{a}$. Consequently, in the later stages of blending, the interpolated $L_{a}$ increasingly resembles $L_{b}$, thus maintaining consistency over time. The average of all previous $L_{a}$ values contributes to the incorporation of general features, particularly during the middle stages of blending, due to the coefficient $p\cdot(1-p)$. Regarding $L^{(t,k)}_{a}$, it gradually declines in the final stages, thereby reducing the proportion of the original features. Through this approach, the auxiliary latents are updated by incorporating both the preceding auxiliary latents and the main blending latent.
Table 1: Quantitative comparisons with other methods. Our method outperforms all other methods across all metrics, including main blending effect (CLIP-BS) , blending objects detection metric DINO-BS, image quality (CLIP-IQA), and better human preference (HPS).
Methods CLIP-BS($\uparrow$) DINO-BS($\uparrow$) CLIP-IQA($\uparrow$) HPS($\uparrow$)
MagicMix (Liew et al., 2022) 8.3063$\pm$ 2.4757 0.2485$\pm$ 0.1575 0.4435$\pm$ 0.0980 0.2710$\pm$ 0.0290
Composable Diffusion (Liu et al., 2022) 6.1374$\pm$ 1.9454 0.2441$\pm$ 0.1655 0.4270$\pm$ 0.1082 0.2903$\pm$ 0.0277
unCLIP-T (Wang et al., 2023) 8.7433$\pm$ 3.1892 0.2214$\pm$ 0.1689 0.4436$\pm$ 0.1073 0.2384$\pm$ 0.0285
unCLIP-U (Wang et al., 2023) 8.7346$\pm$ 3.1577 0.2190$\pm$ 0.1703 0.4450$\pm$ 0.1061 0.2385$\pm$ 0.0280
TEXTUAL (Melzi et al., 2023) 7.8102$\pm$ 2.6852 0.2366$\pm$ 0.1537 0.4164$\pm$ 0.1143 0.2399$\pm$ 0.0230
UNET (Olearo et al., 2024) 8.5080$\pm$ 2.7110 0.2481$\pm$ 0.1628 0.3544$\pm$ 0.2785 0.2405$\pm$ 0.0326
FreeBlend(Ours) 9.1555$\pm$ 2.7134 0.2743$\pm$ 0.1586 0.5238$\pm$ 0.0975 0.2932$\pm$ 0.0316
Table 2: Ablation study for our method with different conditions. We find that the unCLIP-U method, which incorporates two embeddings—one for downsampling and the other for upsampling—achieves the best blending results.
TEXTUAL UNET unCLIP-T unCLIP-U CLIP-BS($\uparrow$)
8.5577$\pm$ 2.6566
9.1555$\pm$ 2.7134
7.8102$\pm$ 2.6852
7.9888$\pm$ 2.5823
8.8170$\pm$ 2.7085
8.5080$\pm$ 2.7110
6.9255$\pm$ 2.4354
9.0260$\pm$ 2.7408
Table 3: Ablation study with different interpolation strategy. The results demonstrate that our increase strategy outperforms both the invariant and decline strategies.
Increase Invariant Decline CLIP-BS($\uparrow$) CLIP-IQA($\uparrow$)
9.1555$\pm$ 2.7134 0.5238$\pm$ 0.0975
7.8891$\pm$ 2.8732 0.4861$\pm$ 0.0990
8.5355$\pm$ 3.0644 0.4981$\pm$ 0.0836
Table 4: Ablation study with staged denoising process. The results demonstrate that our staged denoising process effectively enhances both the blending quality and the scores. Both the initialization and refinement stages contribute to the overall improvement.
Initialization Blending Refinement HPS($\uparrow$)
8.9552$\pm$ 2.7259
8.9832$\pm$ 2.8083
9.1313$\pm$ 2.6941
9.1555$\pm$ 2.7134
4 Experiments
4.1 Experimental Settings
Datasets Construction. We develop the Concept Text-Image Blending (CTIB) dataset, which includes 11,400 text-image pairs. These pairs originate from a collection of 380 text pairs distributed across 20 distinct categories, with each category containing 30 images. The categories are thoughtfully chosen to include a diverse array of real-world objects, highlighting the model’s capability to generate across various species and types.
Implementation Details. Our approach is executed with SD-2-$1$https://huggingface.co/stabilityai/stable-diffusion-2-1 for all baseline comparisons. For our suggested technique, we employ SD-2-1-unCLIP [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1-unclip, enabling image-based conditioning as an input. During the creation of image datasets, the prompt used is “a photo of a {class}”. To maintain a uniform resolution throughout the experiments, all images are resized to $768\times 768$ pixels. The experiments are conducted on a node equipped with 4 NVIDIA GeForce RTX A100 GPUs.
Metrics. Given the unique nature of our task compared to other generation tasks, we use four key metrics to evaluate the blending performance: (1) CLIP Blending Similarity (CLIP-BS) is the primary metric, which measures the distance and similarity between the blending results and the original concepts. (2) DINO Blending Similarity (DINO-BS) (Liu et al., 2025) assesses the detection score of blending objects or the combination of one object with features from another. (3) CLIP Image Quality Assessment (CLIP-IQA) (Wang et al., 2023) evaluates the image quality and the degree of match for the blending objects. (4) Human Preference Score (HPS) (Wu et al., 2023) measures human preferences for the blending results based on blending object prompts. Further explanation is provided in Section A.6.
4.2 Qualitative Comparisons
The results of our comparison with three existing typical concept blending algorithms are illustrated in Figure 3. We observe that using interpolated text embeddings tends to favor one category over the other, resulting in a failure to effectively integrate both elements into the final outcome. On the other hand, while Composable Diffusion successfully introduces both categories, it often leads to concepts co-occurring rather than achieving a harmonious blend. This is evident in cases like “car-cat”, where the concepts remain distinct rather than being seamlessly blended. MagicMix heavily depends on the shape similarity between the reference image and the corresponding category. As a result, it struggles when the categories are completely dissimilar, such as in the cases of “dog-neon light” and “orange-robot”. In these instances, the method fails to generate a meaningful blend, as the lack of visual similarity between the categories hinders the blending process. In contrast, our approach effectively merges both categories, producing well-integrated and balanced results. This enables smooth blending even when the categories are highly distinct, such as in “car-cat” or “dog-neon light”.
4.3 Quantitative Comparisons
We evaluate the performance on our CTIB dataset using four related metrics, with the results presented in Table 1. The results demonstrate that our method significantly outperforms all other approaches across all metrics. This indicates that the quality and the blending effect of our synthetic images achieved by our method is the most visually pleasing.
4.4 Ablation Study
Figure 4: Ablation study on the impact of $\gamma$. The results show that, in the first row, better blending is achieved on the left side, while the right side appears more spliced. In the second row, both sides exhibit a more visually appealing effect. The blending process, however, is inherently subjective, and users can adjust the parameter $\gamma$ to tailor the output according to their preferences. By modulating $\gamma$, users control the contribution of each concept, optimizing performance and addressing associated biases.
Figure 5: The results of user study across our method and three representative concept blending methods. The results demonstrate that our approach achieves the highest level of human preference.
Conditions for Denoising Process. Table 2 presents the results of our ablation study under different conditions. We classify the conditions into two main categories: text and image conditions. For text conditions, we use the TEXTUAL and UNET text prompt blending methods. For image conditions, we apply the unCLIP method along with either unCLIP-T (TEXTUAL-like) or unCLIP-U (UNET-like) image blending methods, where this approach treats the images as text conditions. The results demonstrate that the unCLIP-U method yields the best performance compared to other conditions, while text-based conditions tend to reduce the effectiveness of the generation.
Interpolation Strategy. Table 3 illustrates the impact of different interpolation strategies. In the blending stage, we vary the blending ratio between the interpolated auxiliary latents and the blending latent. From this table, we observe that the “increase” strategy yields a higher CLIP-BS and CLIP-IQA score. As $p$ increases, the blending latent gains more weight relative to the auxiliary latents, resulting in the final output exhibiting features of both kinds. However, the “invariant” method leads to a rigid adjustment, yielding the lowest performance. On the other hand, the “decline” strategy prioritizes the auxiliary latent space and avoids blending, which clearly results in ineffective blending.
Mitigation of bias. We introduce an additional parameter, $\gamma$, to reduce the bias toward specific categories in the generated output image. As shown in Figure 4, we vary $\gamma$ within the range $[0.5,1.5]$. Our observations show that as $\gamma$ moves from 0 to 2, the generated image gradually transitions from resembling reference 1 to favoring reference 2. This provides us with the flexibility to adjust the interpolation strength, allowing for precise control over the image’s characteristics and enabling the achievement of optimal results.
Staged Denoising Process. We validate the impact of our proposed staged denoising process in Table 4. The results demonstrate that both the initialization and refinement stages enhance the stability and quality of the generation. This highlights the effectiveness of the initialization stage in shaping the structure of the initial latents, while the refinement stage successfully adds finer details to the images.
4.5 User Study
We also conduct a user study to evaluate human preference on the blending results with images generated by different methods. Our survey comprised 6 sets of blending pairs. With the active participation of 50 volunteers, we successfully collected a total of 300 votes. The results, as presented in Figure 5, clearly demonstrate the dominance of our method. It received 192 votes, accounting for 64.00% of the total, far surpassing its competitors. Composable Diffusion obtained 63 votes (21.00%), MagicMix received 39 votes (13.00%), and Textual managed only 6 votes (2.00%). These findings not only affirm the effectiveness of our approach but also provide valuable insights into user preferences in the context of concept blending.
5 Conclusion
Blending multiple concepts into a single object based on diffusion models is interesting and valuable in 2D content creation but rather challenging, as it involves controlling the diffusion model to generate previously completely unseen objects. In this paper, we introduce FreeBlend, a training-free method to blend concepts with stark inherent semantic dissimilarities. At its core, FreeBlend employs the unCLIP model with images as conditions and utilizes a novel proposed interpolation strategy with feedback. Extensive qualitative and quantitative studies demonstrate that our method effectively blends disparate concepts, significantly surpassing the prior SoTA counterparts. Moving forward, exploring alternative ways to address the preference dynamics between different concepts may further improve the performance.
Impact Statement
This paper presents a contribution aimed at advancing the field of concept blending. We explicitly state that our work is intended for academic research and commercial applications, provided such uses are authorized by the author. However, we strongly emphasize that we do not condone, nor do we support, the use of our research to generate harmful, unethical, or unlawful content. The generation of unsafe or offensive images that violate ethical standards, legal regulations, or societal norms is strictly prohibited. As researchers, we believe it is our responsibility to ensure that our work is used for the betterment of society and in a manner that aligns with fundamental ethical principles.
Appendix A Appendix
A.1 T-SNE Analysis
In this section, we explore the CLIP embedding space of blended images. Given the high dimensionality of the embeddings, we apply t-SNE (Van der Maaten & Hinton, 2008) for dimensionality reduction and visualize the embeddings in 2D to illustrate their spatial relationships.
Figure 6: Visualization of the CLIP feature distribution for various image categories after dimensionality reduction via t-SNE.
Figure 7: Visualization of the CLIP feature distribution for original class images and blending class images after dimensionality reduction via t-SNE.
In Figure 7, it is evident that semantically related categories tend to be positioned closer to each other, while unrelated categories are spaced farther apart. In Figure 7, we observe that the blending images exhibit a strong relationship with the original classes. Notably, these blending images are generally biased towards one of the original classes, which reflects the inherent bias in the concept blending task. Specifically, we notice that the blending images are not located at the midpoint between the original classes. We hypothesize that this observation can be attributed to the following factors: a) t-SNE is a non-linear dimensionality reduction technique that aims to map similar points in high-dimensional space to nearby points in low-dimensional space, while preserving local structures rather than global relationships. Consequently, the position of points in the t-SNE plot may not necessarily correspond to the relative position of categories in the original feature space. b) The feature vector of the blended class may not be positioned directly between the two original classes. This could be due to: i) The blended class containing a mixture of features from both original classes, with the resulting feature vector being closer to one of the original classes in the feature space. ii) The blended feature may have a direction in the feature space that is relatively distant from both original classes. In certain cases, this could cause the blended class’s feature vector to extend further from the center, leading to its placement behind the original classes in the t-SNE plot, rather than being positioned between them.
A.2 Multiple Concepts Blending.
Our method is not limited to cases involving only two prompts. In Figure 8, we present examples where three or more prompts are incorporated. We utilize the TEXTUAL method to combine three prompts, as the UNET method is limited to handling only two prompts due to the constraints of the U-Net architecture. The blending effect is somewhat chaotic and abstract due to the significant differences and wide conceptual gap among the three prompts. Nevertheless, we managed to achieve relevant results and curated a selection of visually appealing outcomes to present.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
watermelon pineapple tree blending result epiphyllum violet rose blending result
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
cat dog cactus blending result rainbow neon dog blending result
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
bee supercar yellow background blending result rainbow neon dog blending result
Figure 8: Multiple concepts blending results.
Figure 9: Ablation study of the feedback-driven mechanism. Removing the feedback module results in overlapping images, which disrupts the blending process and prevents successful integration.
Figure 10: Visualization of our staged denoising process. From the image, we can observe the effects of the three stages. From the start to $t_{s}$, the image establishes the basic layout of the blending. During the blending stage, both the blending latent and the auxiliary latents are iteratively updated. Finally, in the refinement stage, the quality of the blended image is significantly improved.
Figure 11: Adapting our method to style transfer task.
Figure 12: Screenshot of user study.
A.3 More Ablation Study
Feedback-Driven Mechanism. We conduct an ablation study in Table 5 to validate the effect of the feedback-driven mechanism. The results demonstrate the impact of our feedback-driven mechanism in blending different concepts. The DINO-BS score, however, is not optimal, which may be due to the surprising overlap of blended objects, thus boosting the score. We also present an intuitive visual comparison in Figure 9, where our effect is clearly significant.
Table 5: Ablation study with our feedback mechanism.
Feedback CLIP-BS($\uparrow$) DINO-BS($\uparrow$) CLIP-IQA($\uparrow$) HPS($\uparrow$)
8.9211$\pm$ 3.0076 0.2860$\pm$ 0.1663 0.4874$\pm$ 0.1035 0.2866$\pm$ 0.0305
9.1555$\pm$ 2.7134 0.2743$\pm$ 0.1586 0.5238$\pm$ 0.0975 0.2932$\pm$ 0.0316
Visualization of Staged Denoising Process. Here, we provide a more detailed description of the staged denoising process employed in our method. To facilitate understanding, we visualize the intermediate latent representations, including the blending latent $L_{b}$ and the auxiliary latents $L_{a}^{(t,n)}$ ($n=0,1,2,...,N$) at various stages of the process. The results are presented in Figure 10. Specifically, we focus on timesteps 3 and 22, as they correspond to the initial and final stages where our method begins to exhibit its effects. These visualizations demonstrate how the latents evolve over time, gradually converging to a consistent structure. Through the feedback interpolation module, the blending latent and auxiliary latents seamlessly merge, ultimately resulting in a coherent output image.
A.4 Details of Dataset Construction
Table 6: Categories of Concepts. Categories of these selected concepts for blending are designed to encompass the primary objects commonly found in the real world.
Category Names
Transports airplane, bicycle, car, ship
Animals bird, cat, cattle, dog, fish, squirrel, shark
Common objects chair, desk, phone, robot, lamp
Nature rose, lavender, lemon, water
The designed category list can be divided into four distinct groups based on their characteristics, as shown in Table 6. This categorization provides a clear organization of items based on their respective categories and reflects common concepts encountered in daily life.
A.5 Style Transfer Task
Figure 11 illustrates the impressive visual performance of our method on the style transfer task. Given a reference image and a content image, our training-free approach effortlessly transfers the content into the style of the reference, demonstrating both the adaptability and compatibility of our method for this task.
A.6 Details of Metrics
We calculated the corresponding scores for each image within a category, then computed the average score for that category, and finally determined the overall average score for the entire dataset.
CLIP-BS. For class $A$ and class $B$, we calculate the CLIP blending score using the following equation:
$\text{Score}_{\text{CLIP-BS}}=\sum_{i=1}^{n}\left[\text{Score}(\text{Image}_{\text{blending}},\text{Prompt}_{i})-\text{Score}(\text{Image}_{\text{original}_{i}},\text{Prompt}_{i})\right],$ (12)
where $\text{Score}(\cdot,\cdot)$ is the cosine similarity score computed by the CLIP model using the input image and prompt, $\text{Image}_{\text{blending}}$ represents the generated blended images, $\text{Image}_{\text{original}_{i}}$ denotes the original images of class $i$, and $n$ is the number of concepts blended. $\text{Prompt}_{i}$ refers to the prompt for class $i$, formatted as “a photo of a {class}” to minimize noise in the image generation process. This metric can also be interpreted as the high-dimensional distance between the blended concepts and the original ones. It represents the total distance between the blended concept and each of the original concepts. A larger score generally indicates the dimensional center of the original concepts.
DINO-BS (Liu et al., 2025). Grounding DINO is an open-set object detection model designed to improve the ability to interpret human language inputs through tight modality fusion and large-scale grounded pre-training. The model excels at detecting unseen objects and can perform effective zero-shot transfer learning, even for categories not encountered during training, by leveraging language prompts. To guide the model in detecting blended objects, we designed two specific prompts: “a {concept_a} with blending features from {concept_b}” and “a {concept_b} with blending features from {concept_a}”. These prompts help the model appropriately recognize and identify the blended objects.
CLIP-IQA (Wang et al., 2023). This metric employs a thoughtfully crafted prompt pairing strategy to reduce linguistic ambiguity and effectively harness CLIP’s pretrained knowledge for assessing visual perception. In our task, we design the specific prompt pairs like (“mixed”, “dull”), (“blending features from two different objects”, “natural object from one object”).
HPS (Wu et al., 2023). This metric more accurately predicts human preferences for generated images. The designed prompt is “a photo of a blended object combining mixed features from {concept_a} and {concept_b}”.
A.7 Baseline Methods
In this section, we describe the baseline methods used in our study:
  • MagicMix (Liew et al., 2022). The original MagicMix method modifies the input image using a text prompt from a different class to transfer the image’s characteristics. In our implementation, we first generate the original image using a class text prompt with Stable Diffusion then pass it into the original MagicMix pipeline.
  • Composable Diffusion (Liu et al., 2022). This method decomposes the text description into components, each processed by a different encoder to produce a latent vector capturing its semantic information. Multiple diffusion models are then used to independently generate images for each component. The Conjunction operation combines these components into a single image. This approach excels in zero-shot compositional generation, allowing the model to create novel combinations but not blending of objects and their relationships, even without prior training on such combinations.
  • unCLIP-T (Wang et al., 2023). The unCLIP model leverages images as conditioning inputs, replacing traditional text prompts, by training several linear layer adapters. In this implementation, image embeddings are treated similarly to text embeddings, with their average embedding being computed to serve as the conditions.
  • unCLIP-U (Wang et al., 2023). This implementation is akin to unCLIP-T, with the key difference being that the image embeddings are integrated using the UNET method.
  • TEXTUAL (Melzi et al., 2023). This method encodes the text prompts into text embeddings and computes their average embedding to input into Stable Diffusion for generating blended images. However, it is unstable within the embedding space and constrained by the semantic scope of the concepts involved.
  • UNET (Olearo et al., 2024). This method uses the first text embedding for the conditions of downsampling of U-Net (Ronneberger et al., 2015) and the second one for the upsampling of the U-Net.
A.8 User Study
The screenshots in Figure 12 display our questionnaires used to assess the validation of different methods.
A.9 More Analysis of Blending Results
Blending is a multifaceted process that occurs across various domains, each exhibiting distinct characteristics and behaviors. In the study of embedding blending, it is evident that different categories impose constraints on the resulting blends.
A.9.1 Types of Blending Concept Pairs
Biological Blending. This type of blending involves the blending of entities from the natural world. For example, when animals with similar characteristics, such as cats and dogs, are blended, the resulting output tends to exhibit a high degree of coherence and natural harmony. In contrast, blending animals with plants often leads to situations where the plant components are relegated to background elements, rather than being seamlessly integrated into the final composition. This phenomenon suggests that intra-category blending (e.g., animal-animal or plant-plant) typically yields more naturalistic and fluid results, whereas inter-category blending may result in less cohesive outputs.
Decorative Object-Feature Blending. This form of blending involves the decoration of objects with specific attributes or features. For instance, blending a car with a rose may yield an aesthetic result in the form of a floral-patterned vehicle. Similarly, blending a tomato with a bicycle could result in a red-painted bicycle. In contrast, combining a lily with a bicycle may produce a bicycle adorned with floral elements or set against a backdrop evoking a garden. This type of blending is particularly notable for its capacity to generate creative reinterpretations of objects, often yielding visually appealing and conceptually intriguing combinations.
Material and Pattern Blending. This type refers to the integration of materials—such as stripes, textures, or other patterns—into various objects. The primary characteristic of this form of blending lies in the inherent compatibility of materials with a wide range of object types, thereby facilitating smoother transitions and more seamless integrations across different categories. This suggests that materials have universal applicability, enabling them to blend effectively with diverse objects regardless of their inherent features.
Inherently Unrelated Blending. This form of blending presents significant challenges. When attempting to blend entities from vastly different categories, such as a phone with a dinosaur or an airplane with a dog, the resulting combinations often appear disjointed or nonsensical. This highlights the difficulty in merging categories that exhibit substantial differences in their fundamental characteristics, suggesting that certain combinations may be inherently incompatible.
Broad and General Blending. This type can also complicate the blending process. For example, blending a concept as broad as ”human” tends to result in less coherent outputs, likely due to the wide range of attributes and features that fall under such a generalized category. In contrast, simple, well-defined descriptors such as ”black” tend to blend more effectively, as they are associated with clear, unambiguous characteristics that leave little room for interpretive variation.
Contextual and Background Blending. Context plays a critical role in the outcome of certain blending processes. Many concepts inherently carry with them specific contextual or environmental associations. For example, when the concept of a ”squirrel” is blended, it is often accompanied by a contextual backdrop of trees or forests. These environmental associations are embedded within the very concept of the entity itself, influencing the final blending output and shaping how the resulting image or combination is perceived.
A.9.2 Factors of Blending Concepts
Subject and Background. The effect of concept blending largely depends on the categories to which the blended concepts belong—whether they are subjects or backgrounds. When two concepts, such as a dog and a cat, both belong to the subject category, they are typically blended into a new subject, or one of them may dominate the resulting blend. In contrast, when the blend involves a subject and a background, such as a sheep and a cactus, the result may depict the sheep in a natural environment with cactuses, or the sheep may take on cactus-like characteristics, such as cactus-inspired wool. It is important to note that the identity of a concept can shift depending on how it is paired. For instance, while a cactus serves as the background concept in the pairing ”cactus, sheep,” it becomes part of the subject pair when blended with a rose, as in ”cactus, rose.” Thus, the role of a concept—whether as a subject or background—can significantly alter its representation in the blending process.
Distance between Concepts. Distance is another crucial factor that determines the effectiveness of the blending. This principle is illustrated through t-SNE analysis in Section A.1, which demonstrates how the proximity of two concepts within an embedding space influences their ability to blend. When two concepts are close to each other in this space, they blend easily, forming a cohesive blend. However, as the distance between the concepts increases, the blending becomes less seamless. For example, concepts that are close in proximity, such as cats and dogs, tend to blend effortlessly, with the final output lying somewhere between both concepts. When concepts are moderately distant, such as a cactus and a rose, partial blending occurs, with the resulting image leaning more towards one concept while still maintaining some elements of the other. As the distance between concepts becomes even greater, such as in the case of an airplane and a bird, the blending becomes limited, and the final result often favors one concept, with the other remaining a secondary influence.
Special Cases in blending. Certain concepts possess intrinsic backgrounds or properties that affect blending. For instance, objects like tables and chairs, which are complementary in nature, are more likely to be presented together within the same space rather than being blended into a single form. Similarly, the blending of elements like water and fire may not result in a fluid combination but instead transform into something entirely different, like stone (result from TEXTUAL method), as the inherent properties of the elements conflict with each other.
A.10 Limitations and Future work
Our method has several limitations that negatively affect the blending results. First, the latent interpolation calculation is unstable during generation, often leading to chaotic or overlapping outcomes. Second, when blending multiple concepts (e.g., three or four), the results are unsatisfactory, as the UNET method is not designed to handle more than two conditions simultaneously. Third, the generation process is prone to fluctuation, lacking consistency and stability. Fourth, the Stable Diffusion model exhibits a bias towards different concepts, necessitating the adjustment of the parameter $\gamma$, which is cumbersome and inefficient. Therefore, we aim to explore more stable, cost-effective, and training-free generation methods, as well as seek a more structured approach for blending multiple concepts.
A.11 More Visual Results
Figure 13 and Figure 14 present several of our blending results, highlighting the strong capabilities of our method for both concept blending and generation tasks.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
car-brocolli car-corgi car-lemon car-orange car-teddy
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
orange-brocolli orange-cactus orange-dog orange-cat orange-teddy
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
robot-car robot-cat robot-panda robot-plane robot-brocolli
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
teddy-cat teddy-dog teddy-parrot teddy-neon teddy-penguin
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
eagle-cat eagle-lemon eagle-orange eagle-parrot eagle-penguin
Figure 13: More blending results. Given two concepts, our approach can blend them into a novel object that the model has never seen before, generating high-quality images.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
knight-lemon knight-lion knight-plane knight-panda knight-robot
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
pineapple-blossom pineapple-building pineapple-car pineapple-cat pineapple-train
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
clock-cactus clock-eagle clock-helicopter clock-ice cream clock-lemon
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
ice cream-cat ice cream-lemon ice cream-pineapple ice cream-sakura ice cream-strawberry
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
burger-coca cola burger-dog burger-ice cream burger-robot burger-strawberry
Figure 14: More blending results. Given two concepts, our approach can blend them into a novel object that the model has never seen before, generating high-quality images.